OpenStudy (anonymous):

what is the greatest prime you must consider to test wheather 5084 is prime?? the greatest prime to consider is....

OpenStudy (anonymous):

whatever prime is just below \[\sqrt{5084}\]

OpenStudy (anonymous):

i dont get how to solve that though

OpenStudy (anonymous):

a calculator gives me 71. something, so i would say up until 71

OpenStudy (anonymous):

thank you.. how do i find squar root?

OpenStudy (anonymous):

which is a prime for sure

OpenStudy (anonymous):

how do i do it on a calculator

OpenStudy (anonymous):

actually it is fairly clear that 5084 is divisible by 4. if you divide you get 1271 so i guess i was wrong. you only need to check up to what ever is just below \[\sqrt{1271}\]

OpenStudy (anonymous):

71 was the right answer though

OpenStudy (anonymous):

well if it is on line i guess you can say that. fine

OpenStudy (anonymous):

It seems like, you have to test up to sqrt(n) where n is the number you are testing. So, sqrt(5084) is right.

OpenStudy (anonymous):

You really don't need a calculator in this case. You can tell that 70*70 = 4900. So, it is higher than 70. Then do 71*71 and see how close it gets to 5084. That should be your answer. Of course, calculators are easy to use.

OpenStudy (anonymous):

well only it it is not entirely obvious from your eyeballs that 5084 is divisible by 4.

OpenStudy (anonymous):

The fact that it is divisible by 2 makes it a not-prime number. It does not matter that it is divisible by 4. The largest number you can possibly test is sqrt(n). So, 71 in this case.